OpenStudy (anonymous):

Which ordered pair is a solution of the equation 5x+y=-23? Answer Choices: A. (-3,-4) B.(-5,-3) C.(-4,-3) D.(-3,-5 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (mathstudent55):

I'll show you how you can test the ordered pair (-5, -6) which is NOT one of your choices. Then you can use the same technique for your choices. Is (-5, -6) a solution of the equation 5x + y = -23? Plug in -5 for x and -6 for y in the equation: 5(-5) + (-6) = -23 -25 - 6 = -23 -31 = -23 This is a false statement, so (-5, -6) is not a solution. If you get a true statement, then the ordered pair is a solution.